6. If 6 men can do a piece of work in 14 days, how many men are needed to do the work in 21 days?
a. 3 b. 9 c. 5 d. 4 e. none
The answer on the answer key is D, but I have no clue how to get the same. Please help!

Answer by venugopalramana(3286)

If 6 men can do a piece of work in 14 days, how many men are needed to do the work in 21 days?
there is a simple way to think and do these type of problems.first tabulate as follows.header row with headings.1st.row with given data..2nd row with data to be calculated for ..
.....men....work....days
..... 6..... 1......14
..... ?..... 1......21
Tabulation is clear..now look at the question mark it is in men column.put the number known in that column 6 men first in the numerator (SEE BELOW THIS PARA OF EXPLANATION FOR CALCULATION).go to next column .it is work. read it as follows .6 men required for 1 work (as per first row having given data).how many are needed to do 1 work (as per second row containing question mark) .it is same so multiply now with 1/1 ...now go to next column days.again read it as follows.6 men are needed to do the job in 14 days (as per first row having given data).how many men are needed to do the job in 21 days (as per second row containing question mark)..think there are more days available.so less people will do .so put smaller number of the two 14 in numerator and bigger number 21 in denominator that is 14/21 as the next multiplier. ...like this you can continue for any number of resources,work,etc..every time analyse only one variable for requirement of more or less w.r.t.the data under question mark (here it is men...it could be any thing ..same method and logic will do )..if more is needed put bigger number in numerator and smaller number in denominatot and viceversa...so now the answer is
? = 6 * (1/1) *(14/21) =(6*1*14)/21 =12 is the answer.